Suppose a 2% acid solution is mixed with a 3% acid solution. Find the percent of acid in the following mixture:
a mixture that contains 3 times more 2% acid solution than 3% acid solution.

Answer by stanbon(75887)

Suppose a 2% acid solution is mixed with a 3% acid solution.
Find the percent of acid in the following mixture:
a mixture that contains 3 times more 2% acid solution than 3% acid solution.
----------------------
Let amt of 3% solution be "x".
Then amt. of 2% xolution is "3x"
------------------------------
Quantity of solution = x+3x = 4x
---
Quantity of acid = 0.03x + 0.02(3x) = 0.03 + 0.06x = = 0.09x
-----
Percent of acid = 0.09x/4x = 0.09/4 = 0.0225 = 2.25%
=========================================================
Cheers,
Stan H.